SameikaB:

An expression is shown below: the square root of 18 plus the square root of 2 Which statement is true about the expression? It is rational and equal to 4 times the square root of 2. It is irrational and equal to 4 times the square root of 2. It is rational and equal to 3 times the square root of 2. It is irrational and equal to 3 times the square root of 2.

Luigi0210:

Might be easier if you write it out and simplify: \[\large \sqrt{18}+\sqrt{2}\] \[\large \sqrt{9*2}+\sqrt{2}\] \[\large 3\sqrt{2}+\sqrt{2} =4\sqrt{2}\]